OpenStudy (anonymous):

Can you help me, please : lim(-1)^n * x^n as x approches to +infinity

OpenStudy (anonymous):

@amistre64 ?

OpenStudy (amistre64):

sees loggy to me .... but thats just a thought. what are we trying to determine?

OpenStudy (anonymous):

\[\lim (-1)^{n} * x^{n} as x approaches \to + infinity\]

OpenStudy (amistre64):

as x approaches infinity .... you sure thats not as n to infinity?

OpenStudy (anonymous):

no. it's x

OpenStudy (amistre64):

i dont see it converging for x >= 1

OpenStudy (anonymous):

So ? @amistre64

OpenStudy (amistre64):

im going to go out on a limb here and say ive got no idea what you are trying to ask for ...

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

please wait a moment

OpenStudy (michele_laino):

If I call with a_n this quantity: \[\Large {a_n} = {\left( { - 1} \right)^n}{x^n}\]

OpenStudy (michele_laino):

then we can write: \[\Large {a_n} = {x^n}\] if n is an even number and its limits is +infinity whereas, we have: \[\Large {a_n} = - {x^n}\] if n is an odd number, and its limit is -infinity so what we can conclude?

OpenStudy (michele_laino):

I supposed that x>0, and x is different from zero

OpenStudy (anonymous):

so the result would be.... ?

OpenStudy (michele_laino):

we conclude that the limit value doesn't exist

OpenStudy (anonymous):

ok... what abou this : lim\[\sin ^{2}x+\tan 2x + \sqrt{x}\] as x approaches to pi Is sqrt of pi the result ?

OpenStudy (michele_laino):

we have: sinx--->0, so (sinx)^2--->0*0=0 tan(2x)--->0 and: sqrt(x)--->sqrt(pi) so your answer is right!

OpenStudy (anonymous):

what about \[\lim (\sqrt{2x+1}-\sqrt{4x+1})/\sqrt[3]{x} as x approaches \to 0\]

OpenStudy (michele_laino):

we have to transform your function, as below: \[\Large \begin{gathered} \frac{{\sqrt {2x + 1} - \sqrt {4x + 1} }}{{\sqrt[3]{x}}} \times \frac{{\sqrt {2x + 1} + \sqrt {4x + 1} }}{{\sqrt {2x + 1} + \sqrt {4x + 1} }} = \hfill \\ \hfill \\ = \frac{{2x + 1 - 4x - 1}}{{\sqrt[3]{x}\left( {\sqrt {2x + 1} + \sqrt {4x + 1} } \right)}} = - \frac{{2\sqrt[3]{{{x^2}}}}}{{\sqrt {2x + 1} + \sqrt {4x + 1} }} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

now, in order to compute your limit, you have to set x=0, into the last expression

OpenStudy (anonymous):

Hmmm...right ! and the last one please \[\lim (x^{n}+1)/x^{2} \] as x approaches to +infinity and n is in N

OpenStudy (michele_laino):

we can write your expression as follows: \[\Large {x^{n - 2}} + \frac{1}{{{x^2}}}\]

OpenStudy (michele_laino):

now, if n<2, then n-2=-m<0, so your limit is equal to: the limit of this function: \[\frac{1}{{{x^m}}} + \frac{1}{{{x^2}}} \to ...?\]

OpenStudy (michele_laino):

please continue

OpenStudy (michele_laino):

what value do you get?

OpenStudy (anonymous):

0?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

now, if n=2, then n-2=0, so your limit is the limit of this function: \[\frac{1}{{{x^2}}} \to ...?\]

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

0 :D

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

finally, if n>2, then n-2=q>0, so your limit is the limit of this function: \[{x^q} + \frac{1}{{{x^2}}} \to ...?\]

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

what a minute ,, we have x^q ... that confuses me

OpenStudy (michele_laino):

yes! and q>0

OpenStudy (anonymous):

infinity + 0 ?!

OpenStudy (michele_laino):

yes! and what is: +infinity+0=...?

OpenStudy (anonymous):

infinity ?

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

so we have discovered that: if n=2 or n<2, then your limit is zero, whereas if n>2, then your limit is +infinity

OpenStudy (anonymous):

If you could see my face right now, you'd see a look of gratitude :)

OpenStudy (michele_laino):

thanks! :)

OpenStudy (anonymous):

I am indebted to you...Thank you so much :))))))

OpenStudy (michele_laino):

thanks again! :)

OpenStudy (michele_laino):

I'm very sorry, I have made an error: in the case n=2, your limit is the limit of this function: \[\Large {x^0} + \frac{1}{{{x^2}}} = 1 + \frac{1}{{{x^2}}} \to 1 + 0 = 1\]